### v1

# Phase driven exact and unconventional superradiance phase transition in non-Hermitian cascaded Rabi cavities

Abstract
This work reports the phase driven symmetry breaking and exact and unconventional superradiance phase transition in the non-Hermitian cascaded Rabi cavities. The non-Hermiticity is introduced in the coupling phase (denoted by $\varphi$) between the atom and the optical field. The exactness refers to the fact that the superradiance phase boundary is obtained analytically and verified by the observables. The unconventionality is reflected in that when $|\varphi|=\frac{\pi}{4}$ or $|\varphi|=\frac{3\pi}{4}$, the phase boundary is uniquely determined by $\mathcal{J}=\frac{1}{2}$ (where $\mathcal{J}$ is the dimensionless cavity coupling strength) and is independent of the atom-optical field coupling strength $g$. For other $\varphi$, the phase boundary is determined by $\mathcal{J}$ and the dimensionless atom-optical field coupling strength $g$ together. Besides, we find that there are phase driven first-order and second-order superradiance phase transitions, and the quantum criticality for the second-order superradiance phase transition is studied. In addition, the experimental feasibility is discussed. This work will stimulate the studies of non-Hermitian superradiance quantum phase transitions and their experimental realizations, as well as the underlying universality class of phase transitions.
pacs:
64.60.-i, 05.70.Fh, 64.60.Fr
I Introduction
The quantum Rabi model (QRM) named as Rabi Rabi ; book describes the interaction of a two-level atom and a single-model optical field and can be viewed as the minimal version of the Dicke model Dicke . In the past years, there are efforts in finding the exactly analytically solutions of the QRM am_1 ; am_3 ; am_2 ; am_4 ; am_5 ; am_6 ; am_7 ; am_8 . Employing the Bargmann space representation bs , Braak obtained the transcendental function, namely the G-function, and found that the poles of the G-function were the exact solutions of QRM es1 . Quickly, Chen et al. obtained the exact solutions of a two-photon QRM by combining the Bogoliubov transformation and the G-function es2 . In the decade, the two methods have been used for calculating the exact solutions of the generalized QRMs ext_1 ; ext_2 ; ext_3 ; ext_4 ; ext_5 ; ext_6 ; ext_7 ; ext_8 ; ext_9 ; ext_10 ; ext_11 . It was found that the symmetry breaking will lead to the topological transition in the QRM tp1 and the generalized QRM tp2 .
Accompanying by the search for the exact solutions, there are interests in studying the quantum criticality of the QRMs or the generalized QRMs HP ; ps1 ; ps2 ; R_d1 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 ; zhangyy_2 . Accompanied by the quantum phase transition, i.e., the superradiance phase transition (SPT), there will occur macroscopic photon excitations R_d1 . In 2015, Hwang et al. studied the universal dynamics and the critical behaviors of the QRM and the exact phase boundary of SPT was obtained by an effective low-energy approximation HP . They found that the critical exponent of SPT $\nu$ is $\nu=\frac{3}{2}$ HP . By analyzing a general scaling function, Liu et al. found that the $\nu$ in the anisotropic QRM sill satisfies $\nu=\frac{3}{2}$ and the phase boundary of SPT is obtained by a semi-class method ps2 . Quickly, Wei et al. numerically verified the exact $\nu$ by analyzing the critical behaviors of the finite frequency fidelity susceptibility qc_1 . Subsequent researches show that for a multi-cavity system with two qc_2 ; qc_3 or three zhangyy_2 cascaded QRMs, the crtical exponent is still equal to that of a single Rabi model, but for a three-level QRM, there occurs new critical exponent with $\nu=1$zhangyy_3 . We notice that the SPTs in the QRMs more or less depend on the atom-optical field coupling strength. It motivate us whether there exist an unconventional SPT which is independent of the atom-optical field coupling.
Recently, the combination of the non-Hermiticity and the QRMs triggered intriguing physics. Cao et al. constructed a generalized non-Hermitian integrable QRM. They obtained the exact energy spectrum and eigenstates of this model by employing the Bethe ansatz es_NH_1 . The work provides a generalized strategy to engineer integrable non-Hermitian spin-boson models. By introducing the parity-time ($\mathcal{PT}$) symmetric gain and loss in the semi-classical Rabi model, Lee et al. PT_Rabi_1 found that there was $\mathcal{PT}$ symmetry breaking PT ; PT1 accompanied by the real-complex transition in energy. With the exact solutions of the $PT$-symmetric Rabi models, one can locate the $PT$-symmetry breaking point, i.e., the exceptional point (EP) EP . Xie et al. investigated the Floquet $\mathcal{PT}$-symmetric semi-classical Rabi model es_NH_2 . Under the multiple-photon resonance condition, they obtained the positions of the EPs from the exact Floquet spectrum. Employing the adiabatic approximation, Lu et al. obtained the exact energy spectrum of the $\mathcal{PT}$-symmetric QRM, by which the EPs were located es_NH_3 . With the emergence of EP, observables present discontinuous behaviors. By the Bogoliubov transformation and G-function, Li et al. solved the exact energy spectra for a one-photon and two-photon $\mathcal{PT}$-symmetric QRM, and the zeros of G-function located the EPs es_NH_4 . These works provide an inspiration for us to study whether there will be a completely real energy spectrum in a generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry. In addition, we will study the quantum criticality of the model and discuss the similarity and difference with Hermitian cases.
The paper is organized as follows. In Sec. II we introduce the generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry. In Sec. III, we obtain the exact phase boundary of the model. In Sec. IV, we study the observables which signal the SPT. In Sec. V, we discuss the quantum criticality of the SPT. A experimental feasibility is discussed in Sec. VI, and a summary is presented in Sec. VII.
II Two coupled non-Hermtian Rabi cavities
The generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry is constructed by two cascaded non-Hermitian Rabi cavities, whose Hamiltonian ($\hbar=1$) is given as
$H=\sum_{j=1,2}h_{j}+J(a^{{\dagger}}_{1}+a_{1})(a^{{\dagger}}_{2}+a_{2}),$ (1)
where $h_{j}$ describes non-Hermitian Rabi cavity $h_{j}=\omega_{p}a^{{\dagger}}_{j}a_{j}+\frac{\omega_{a}}{2}\sigma^{j}_{z}-\lambda e^{i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}$ with $j$ the cavity index and $a^{{\dagger}}_{j}$ ($a_{j}$) the creation (annihilation) of photons. $\omega_{p}$ is the optical field frequency, $\omega_{a}$ is the atomic frequency, and $\lambda e^{i\varphi}$ is the coupling strength between the atom and optical field with $\lambda$ the weight of strength and $\varphi$ the coupling phase. The non-Hermiticity of the system requires $\varphi\in(-\pi,0)\cup(0,\pi)$.
Figure 1: (Color online) Ground-state phase diagram in the $\mathcal{J}$-$g$ parameter space. (a) $\varphi=\frac{\pi}{2}$. (b) $\varphi=\frac{\pi}{4}$. The blue solid lines denote the phase boundary, satisfying $1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)=2\mathcal{J}$.
For a single non-Hermitian Rabi cavity, we argue that there is no parity-time ($\mathcal{PT}$) symmetry unless the coupling phase $\varphi=\pm\frac{\pi}{2}$. The time-reversal operator $\mathcal{T}$ dominates the complex conjugate operation, resulting in $\mathcal{T}a^{{\dagger}}_{j}(a_{j})\mathcal{T}$=$a^{{\dagger}}_{j}(a_{j})$. The parity operator $\mathcal{P}$ dominates the inversion operation, i.e., $\mathcal{P}\hat{x}_{j}\mathcal{P}$= $-\hat{x}_{j}$, and $\mathcal{P}\hat{p}_{j}\mathcal{P}$= $-\hat{p}_{j}$ PT . Thus, we have
$\mathcal{PT}h_{j}\mathcal{TP}$ $=\omega_{p}a^{{\dagger}}_{j}a_{j}-\mathcal{P}\left[\lambda e^{-i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}\right]\mathcal{P}+\frac{\omega_{a}}{2}\sigma^{j}_{z}$ (2)
$=\omega_{p}a^{{\dagger}}_{j}a_{j}+\lambda e^{-i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}+\frac{\omega_{a}}{2}\sigma^{j}_{z}.$
Only if $\varphi=\pm\frac{\pi}{2}$, the single non-Hermitian Rabi cavity is $\mathcal{PT}$-symmetric (see another prove in Ref. es_NH_4 ). For other values, it is not $\mathcal{PT}$-symmetric.
For the cascaded non-Hermitian QRM, we demonstrate that this system is completely not $\mathcal{PT}$-symmetric. $\mathcal{P}$ dominates a mirror operation between the two cavities, i.e., $\mathcal{P}a^{{\dagger}}_{j}(a_{j})\mathcal{P}=a^{{\dagger}}_{j^{\prime}}(a_{j^{\prime}})$, $\mathcal{P}\sigma^{j}_{z}(\sigma^{j}_{x})\mathcal{P}=\sigma^{j^{\prime}}_{z}(\sigma^{j^{\prime}}_{x})$ (here $j\neq j^{\prime}$), and $\mathcal{T}$ leads to $\mathcal{T}i\mathcal{T}=-i$. Performing $\mathcal{PT}$ on the cavity coupling term, we have
$\mathcal{PT}J\left(a^{{\dagger}}_{1}+a_{1}\right)\left(a^{{\dagger}}_{2}+a_{2}\right)\mathcal{TP}=J\left(a^{{\dagger}}_{1}+a_{1}\right)\left(a^{{\dagger}}_{2}+a_{2}\right).$ (3)
Performing $\mathcal{PT}$ on $h_{j}$, we have
$\mathcal{PT}h_{j}\mathcal{TP}$ $=\omega_{p}a^{{\dagger}}_{j^{\prime}}a_{j^{\prime}}-\lambda e^{-i\varphi}\left(a_{j^{\prime}}+a^{{\dagger}}_{j^{\prime}}\right)\sigma^{j^{\prime}}_{x}+\frac{\omega_{a}}{2}\sigma^{j^{\prime}}_{z}$ (4)
$\neq h_{j^{\prime}}.$
Therefore, here the non-Hermitian cascaded QRM is not $\mathcal{PT}$-symmetric.
III Exact and unconventional superradiance phase transition
To analytically solve the phase boundary of the ground-state superradiance phase transition, we introduce the dimensionless parameters $\eta=\frac{\omega_{a}}{\omega_{p}}$, $g=\frac{2\lambda}{\sqrt{\omega_{a}\omega_{p}}}$, and $\mathcal{J}=\frac{J}{\omega_{p}}$. Then, the Hamiltonian in Eq. (1) can be rewritten as
$\frac{H}{\omega_{p}}=\sum_{j=1,2}\frac{h_{j}}{\omega_{p}}+\mathcal{J}\left(a^{{\dagger}}_{1}+a_{1}\right)\left(a^{{\dagger}}_{2}+a_{2}\right),$ (5)
in which
$\frac{h_{j}}{\omega_{p}}=a^{{\dagger}}_{j}a_{j}+\frac{\eta}{2}\sigma^{j}_{z}-\frac{g\sqrt{\eta}}{2}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}.$ (6)
Performing the following transformation
$x_{j}$ $=\left(a^{{\dagger}}_{j}+a_{j}\right)/\sqrt{2\eta},$ (7)
$y_{j}$ $=i\left(a^{{\dagger}}_{j}-a_{j}\right)\sqrt{\eta/2},$
and dividing both sides of Eq. (5) by $\eta$, we obtain
$\mathcal{H}=\frac{H}{\omega_{p}\eta}\equiv\frac{1}{2}\sum_{j=1,2}\left(x^{2}_{j}+\frac{y^{2}_{j}}{\eta^{2}}+\sigma^{j}_{z}-\sqrt{2}g\sigma^{j}_{x}x_{j}\right)+2\mathcal{J}x_{1}x_{2}.$ (8)
As studied in Refs. ps1 ; ps2 , the SPT appears in the under the infinite $\eta$ limit. Under this condition, we can obtain the analytical phase boundary of SPT. When $\eta\rightarrow\infty$, $\mathcal{H}$ reduces to
$\mathcal{H}=\frac{1}{2}\sum_{j=1,2}\left[x^{2}_{j}+\sigma^{j}_{z}-\sqrt{2}g\sigma^{j}_{x}x_{j}\right]+2\mathcal{J}x_{1}x_{2}.$ (9)
By diagonalizing $\mathcal{H}$, we obtain the energies $E_{\pm}(x_{1},x_{2})$ as
$E_{\pm}(x_{1},x_{2})=\frac{1}{2}\sum_{j=1,2}\left(x^{2}_{j}\pm\sqrt{1+2g^{2}x^{2}_{j}}\right)+2\mathcal{J}x_{1}x_{2}.$ (10)
Performing the Taylor expansion on $E_{\pm}(x_{1},x_{2})$, and truncating them to the quadratic term $\mathcal{O}(x^{4}_{j})$, the low-energy branch $E_{-}(x_{1},x_{2})$ (contains the ground-state information) becomes
$E_{-}(x_{1},x_{2})=\frac{1}{2}\psi^{T}\Lambda\psi,$ (11)
in which $\psi=(x_{1},x_{2})^{T}$ and
$\Lambda=\begin{pmatrix}1-g^{2}e^{i2\varphi}&2\mathcal{J}\\ 2\mathcal{J}&1-g^{2}e^{i2\varphi}\end{pmatrix}.$ (12)
Evidently, the matrix $\Lambda$ consists of the real part ${\rm Re}(\Lambda)$ and the pure imaginary part ${\rm Im}(\Lambda)$ (which can be considered as the background loss), i.e., $\Lambda={\rm Re}(\Lambda)+{\rm Im}(\Lambda)$, with
${\rm Re}(\Lambda)$ $=\left[1-g^{2}(\cos^{2}\varphi-\sin^{2}\varphi)\right]\sigma_{0}+2\mathcal{J}\sigma_{x},$ (13)
${\rm Im}(\Lambda)$ $=-i2g^{2}\cos\varphi\sin\varphi\sigma_{0}.$
Figure 2: (Color online) The real part of the ground-state energy ${\rm Re}(E_{0})/\eta$ as a function of $\mathcal{J}$ under different $\eta$ with $g=0.7$ and $\phi=\frac{\pi}{2}$ in (a) and with $g=0.4$ and $\varphi=\frac{\pi}{4}$ in (b). (c) The second-order derivative of ${\rm Re}(E_{0})/\eta$ for $g=0.7$ and $\phi=\frac{\pi}{2}$. (d) The first-order derivative of ${\rm Re}(E_{0})/\eta$ for $g=0.4$ and $\phi=\frac{\pi}{4}$. The photonic truncation number is $N=80$.
In fact, the phase boundary of SPT is nothing to do with background loss ${\rm Im}(\Lambda)$, but is determined by ${\rm Re}(\Lambda)$. After diagonalizing ${\rm Re}(\Lambda)$, the eigenvalues ϵ±subscriptitalic-$\epsilon_{\pm}$ of ${\rm Re}(\Lambda)$ is obtained as $\epsilon_{\pm}=1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)\pm 2\mathcal{J}$. When the low-energy branch in ϵ±subscriptitalic-$\epsilon_{\pm}$ equals to zero, i.e, $\epsilon_{-}=0$, we can extract the phase boundary of the superradiance phase transition, and it satisfies $1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)=2\mathcal{J}$. From the expression of the phase boundary, we can see that there are special phases, i.e., $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, at which the phase boundary is fixed at $\mathcal{J}=\frac{1}{2}$, and is independent of the atom-optical field coupling parameter $g$. Such type of SPT independent of atom-optical field coupling is counter-intuitive and has not been reported before, so we name this type of SPT as unconventional SPT. While $\gamma$ is taken at other values, the phase boundary depends on the $g$ and $\mathcal{J}$ together. Thus, we name this atom-optical field coupling dependent SPT as the conventional SPT. Without loss of generality, in the following, we will study the energy, photon population, qubit population, the fidelity and the fidelity susceptibility in the ground state at $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ to characterize the properties of the SPT of the non-Hermitian two-Rabi cavity system. The ground state here refers to the state with the smallest real part of the energy. Intuitively, the phase diagrams for $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ are plotted in Fig. 1(a) and 1(b), respectively. The red regions denote the superradiance phases, separated from the normal phases (yellow regions) by the phase boundaries (blues curves).
Figure 3: (Color online) The ground-state photon population $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ as a function of $\mathcal{J}$. (a) $g=0.5$ and $\phi=\frac{\pi}{2}$; (b) $g=0.7$ and $\varphi=\frac{\pi}{2}$; (c) $g=0.4$ and $\varphi=\frac{\pi}{4}$; (d) $g=0.6$ and $\varphi=\frac{\pi}{4}$. The photonic truncation number is $N=80$ and $\eta=1500$.
IV Signatures of the superradiance phase transition
We detect the signatures and check the phase boundary of the SPT. We start by studying the ground-state energy to reveal the SPT. The energies are sorted by the real parts of the energies, and the ground-state energy means the energy with the lowest real part. Under $\varphi=\frac{\pi}{2}$ and different $\eta$, the real parts of the ground-state energies ${\rm Re}(E_{0})/\eta$ as a function of $\mathcal{J}$ for $g=0.7$ are plotted in Fig. 2(a). The photonic truncation number is taken at $N=80$. As can be seen that with the emergence of SPT, ${\rm Re}(E_{0})/\eta$ for different $\eta$ decrease from $-1$ to lower values. According to the obtained analytical result, the transition point under $\varphi=\frac{\pi}{2}$ and $g=0.7$ should be at $\mathcal{J}=0.745$. Intuitively, the SPT occurs at the transition points predicted by the exactly analytical result (shown with the vertical black line). Meanwhile, we calculate ${\rm Re}(E_{0})/\eta$ in the $\varphi=\frac{\pi}{4}$ case and ${\rm Re}(E_{0})/\eta$-$\mathcal{J}$ relations for $g=0.4$ are plotted in Fig. 2(b). Similar to the $\varphi=\frac{\pi}{2}$, when $\mathcal{J}$ exceeds to the transition point, ${\rm Re}(E_{0})/\eta$ for different $\eta$ in the $\varphi=\frac{\pi}{4}$ case decrease from $-1$ to lower values, signaling the appearance of SPT as well. From the analytical expression of phase boundary, it is known that the transition point under $\varphi=\frac{\pi}{2}$ should be at $J=0.5$ regardless of what value of $g$ is. It is readily seen that for $g=0.4$, the transition points of ${\rm Re}(E_{0})/\eta$ occur at the analytically obtained transition points. Although the ground state energy of the system decreases after the occurrence of conventional and unconventional SPTs, it should be pointed out that they do not belong to the same phase transition type. As shown in Fig. 1(c), the second-order derivative of ${\rm Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{2}$ and $g=0.7$ is not continuous at the the transition point. Figure 1(d) shows that the first-order derivative of ${\rm Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{4}$ and $g=0.4$. The results show that the conventional SPT should be the second-order quantum phase transition, while the unconventional SPT should be the first-order quantum phase transition. In fact, only the $|\varphi|=\frac{\pi}{4}$ and $|\varphi|=\frac{3\pi}{4}$ cases correspond to the second-order quantum phase transition, while the cases with other $\varphi$ correspond to the first-order quantum phase transition. However, we are more concerned with conventional and unconventional superradiance phase transitions in this paper, so the results of other $\varphi$ are not considered
Figure 4: (Color online) The ground-state atom population $\sigma^{1(2)}_{z}$ as a function of $\mathcal{J}$. (a) $g=0.5$ and $\phi=\frac{\pi}{2}$; (b) $g=0.7$ and $\varphi=\frac{\pi}{2}$; (c) $g=0.4$ and $\varphi=\frac{\pi}{4}$; (d) $g=0.6$ and $\varphi=\frac{\pi}{4}$. The photonic truncation number is $N=80$ and $\eta=1500$.
Except for the ground-state energy, the phenomenon of SPT can be readily seen from the ground-state photon population. Under $\varphi=\frac{\pi}{2}$, ground-state photon populations $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ for $g=0.5$ and $g=0.7$ are plotted in Fig. 3(a) and 3(b), respectively. It is seen that there is no photon excitation when $\mathcal{J}$ is smaller than the transition point, but there are macroscopic photon excitations when the cavity coupling parameter $\mathcal{J}$ exceeds the transition point. The phenomenon is similar to the $\varphi=\frac{\pi}{4}$ case. For $g=0.4$ and $g=0.6$, the corresponding ground-state photon populations are plotted in Fig. 3(c) and 3(d), respectively. We can see that there are macroscopic photon excitations when $\mathcal{J}$ is larger than $0.5$, while the photon excitations are evidently suppressed when $J<0.5$.
A reasonable explaination for the macroscopic photon excitations in the suprradiance phase is that the atoms are excited and become active. To check this point, the atom populations $\langle\sigma^{1(2)}_{z}\rangle$ as a function of $\mathcal{J}$ for different $\varphi$ and $g$ are plotted in Fig. 4(a)-4(d). We can see that when $\mathcal{J}$ is smaller than the transition point, the atoms are stable in the ground state, with populations $\langle\sigma^{1(2)}_{z}\rangle=-1$. When $\mathcal{J}$ crosses the transition point, it is seen that the populations of atoms deviate from -1, implying that the atoms are excited and become active. The excited atoms dominate the macroscopic photon excitations in the SPT. Although the photon excitations and the atom populations of the system vary obviously with occurrence of SPT, the two types of SPTs lead to different varying behaviors of observables. From Fig. 3, we can see that when the conventional SPT happens, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ continuously varies to macroscopic value, while the unconventional SPT happens, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ directly jumps to macroscopic values without asymptotic process. Similarly, we can see the obvious difference between the two types of SPTs from the atomic populations in Fig. 4.
Figure 5: (Color online) The excitation gaps ${\rm Re}(\Delta_{E})$ and ${\rm Im}(\Delta_{E})$ as functions of $\mathcal{J}$ under different $\eta$. (a1) and (a2): $g=0.5$ and $\phi=\frac{\pi}{2}$; (b1) and (b2): $g=1$ and $\varphi=\frac{\pi}{2}$; (c1) and (c2): $g=0.4$ and $\varphi=\frac{\pi}{4}$; (d1) and (d2): $g=0.6$ and $\varphi=\frac{\pi}{4}$. The photonic truncation number is $N=36$.
In addition to the photon population and the atom population, the excitation gaps behind the two types of SPTs are different as well. The excitation gaps here contain two meanings. One is the real excitation gap ${\rm Re}(\Delta_{E})/\eta$ between the real parts of the lowest two energies. Another is the imaginary excitation gap ${\rm Im}(\Delta_{E})/\eta$ between the imaginary parts of the lowest two energies. Figures. 5(a1)-5(d1) (Figs. 5(a2)-5(d2)) presents ${\rm Re}(\Delta_{E})/\eta$ (${\rm Im}(\Delta_{E})/\eta$) as a function of $\mathcal{J}$ under different $\eta$. We can see that no matter $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$, the real excitation gap experiences a gap closing process when $\mathcal{J}$ exceeds the transition point. However, for the imaginary excitation gap, there is an obvious difference between the two studied cases. From Figs. 5(a2) and 5(b2), we can see that ${\rm Im}(\Delta_{E})/\eta$ always keeps zero (${\rm Im}(\Delta_{E})/\eta$ far less than $10^{-14}$ can be numerically regarded as zero.) before and after the transition point. From Figs. 5(c2) and 5(d2), we can see that the imaginary excitation gaps experience the gap closing process, similar to the circumstances for the real excitation gaps.
Figure 6: (Color online) The maximal imaginary part of the energy spectrum ${\rm Im}(E)_{max}/\eta$ as functions of $\mathcal{J}$ under different $\eta$. (a) $\varphi=\frac{\pi}{2}$ and $g=0.7$. (b) $\varphi=\frac{\pi}{2}$ and $g=0.8$. (c) $\varphi=\frac{\pi}{4}$ and $g=0.4$. (d) $\varphi=\frac{\pi}{4}$ and $g=0.6$. The photonic truncation number is $N=36$.
It is noted that for a $\mathcal{PT}$-symmetric Rabi system, there exists real-complex transition in energy spectrum PT ; PT1 ; es_NH_2 ; es_NH_3 ; es_NH_4 . Surprisingly, for the non-Hermitian two-Rabi cavity system without $\mathcal{PT}$ symmetry, there exists real energy spectrum as well. Under $\varphi=\frac{\pi}{2}$, the maximal imaginary of the energy spectra for $g=0.7$ and $g=0.8$, i.e., ${\rm Im}(E)_{max}$ are plotted in Figs. 6(a) and 6(b), respectively. The analytically obtained superradiance transition point for $g=0.7$ is $J=0.745$, and that for $g=0.8$ is $J=0.82$. ${\rm Im}(E)_{max}/\eta$ approaches to zero before and after the superradiance transition point implying that the energy spectra under $\varphi=\frac{\pi}{2}$ are real, and there is no real-complex transition in energy spectrum. Compared to the $\varphi=\frac{\pi}{2}$ case, the spectral property of the $\varphi=\frac{\pi}{4}$ is different. Under $\varphi=\frac{\pi}{4}$, ${\rm Im}(E)_{max}/\eta$ for $g=0.4$ and $g=0.6$ are plotted in Figs. 6(c) and 6(d), respectively. Different from the $\varphi=\frac{\pi}{2}$ cases, here the energy spectra are complex because of the nonvanishing ${\rm Im}(E)_{max}/\eta$. In fact, the real and complex characteristics of the energy spectrum can be understood analytically as well. From Eq. (10) and Eq. (13), we know that when $\varphi=\frac{\pi}{2}$, the imaginary part disappears, but when $\varphi=\frac{\pi}{4}$, the imaginary part still exists. The results in Fig. 6 provide numerical validation of the previous analytical derivation. Considering the fact that the $\varphi=\frac{\pi}{2}$ cases have full real energy spectra, we conjecture that here the conventional SPT belongs to the same universality class as the Hermitian Rabi systems HP ; ps2 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 and the critical exponent of the phase transition can be obtained from the critical behavior. However, for the complex energy spectrum case ($\varphi=\frac{\pi}{4}$), we conjecture that there is no critical behavior. Although the existence of quantum criticality of conventional SPT and the absence of quantum criticality of unconventional SPT have been reflected in the photon population and atom population, we still want to further verify them by studying the ground-state fidelity and fidelity susceptibility.
Figure 7: (Color online) Fidelity susceptibility $\chi_{\mathcal{F}}$ as functions of $\mathcal{J}$ under different $\eta$ with $\varphi=\frac{\pi}{2}$ and $g=0.7$ in (a), and with $\varphi=\frac{\pi}{2}$ and $g=0.8$ in (b). The involved parameter is $\delta\mathcal{J}=10^{-5}$. Fidelity $\mathcal{F}$ as functions of $\mathcal{J}$ under different $\eta$ with $\varphi=\frac{\pi}{4}$ and $g=0.4$ in (c), and with $\varphi=\frac{\pi}{4}$ and $g=0.6$ in (c). The involved parameter is $\delta\mathcal{J}=5\times 10^{-4}$. The photonic truncation number in all calculations is $N=80$.
V conventional and unconventional superradiance phase transition
We focus on the quantum phase transition and its critical behavior triggered by the cavity coupling parameter $\mathcal{J}$. The fidelity for the ground state $\mathcal{F}$ is defined as fidelity_0 ; fidelity_1 ; fidelity_2 ; fidelity_3
$\mathcal{F}=|\langle\psi_{0}(\mathcal{J})|\psi_{0}(\mathcal{J}+\delta\mathcal{J})|,$ (14)
which weighs the survival probability of the ground state after the cavity coupling parameter $\mathcal{J}$ from an infinitesimal distance $\delta\mathcal{J}$. Further, according to Refs. FS_1 ; FS_2 ; FS_3 ; FS_4 ; FS_5 , we can obtain the fidelity susceptibility $\chi_{\mathcal{F}}$ as
$\chi_{\mathcal{F}}=\lim_{\delta\mathcal{J}\rightarrow 0}\frac{-2\ln\mathcal{F}(\mathcal{J},\delta\mathcal{J})}{(\delta\mathcal{J})^{2}}.$ (15)
If there exists quantum phase transition, the fidelity will experience a drop at the phase transition point, and the fidelity susceptibility will peak at the phase transition point.
Figure 8: (Color online) (a) Universal finite $\eta$ scaling relation of the maximal fidelity susceptibility $\chi^{max}_{F}$. (b) Data collapse of $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}$. The involved parameters are $\varphi=\frac{\pi}{2},$$g=0.7$, and $\delta J=10^{-5}$.
Under $\varphi=\frac{\pi}{2}$, $\chi_{\mathcal{F}}/\eta$ for $g=0.7$ and $g=0.8$ are plotted in Figs. 7(a) and 7(b), respectively. As seen that with the increase of $\mathcal{J}$, $\chi_{\mathcal{F}}/\eta$ smoothly reaches the peak and then smoothly decreases. Besides, $\chi_{\mathcal{F}}/\eta$ presents obvious critical behaviors, because it is seen that with the increase of $\eta$, the peak of $\chi_{\mathcal{F}}$ at the coupling parameter $\mathcal{J}_{max}$ gradually tends to the exact solution. Moreover, with the increase of $\eta$, the peak of $\chi_{\mathcal{F}}$ get higher. However, for the $\varphi=\frac{\pi}{4}$ cases, there is no critical phenomenon. Under $\varphi=\frac{\pi}{4}$, the fidelity $\mathcal{F}$ for $g=0.4$ and $g=0.6$ are plotted in Figs. 7(c) and 7(d), respectively. Intuitively, $\mathcal{F}$ sharply drop to zero at single parameter $\mathcal{J}$ and there is no asymptotic behavior between the valley of $\mathcal{F}$ (or the peak of $\chi_{\mathcal{F}}$) and $\eta$. In other words, the conventional SPT exists quantum criticality while there is no quantum criticality for the unconventional SPT.
Next, we study the quantum criticality of the conventional SPT ($\varphi=\frac{\pi}{2}$ case). From the $\chi_{\mathcal{F}}$-$\eta$ curves, the maximal fidelity susceptibility $\chi^{max}_{F}$ under a specific $\eta$ is obtained. Then, the universal finite $\eta$ scaling relation of $\chi_{\mathcal{F}}$ for $g=0.7$ is plotted in Fig. 8(a). The black fitting curve satisfies $\ln\chi_{\mathcal{F}}=(-1.363\pm 0.004)\ln\eta+2.388$. It implies the scaling exponent $\mu$ of $\chi_{\mathcal{F}}$ is $\mu=-1.363\pm 0.004$. As introduced in Ref. qc_1 ; RMP2011 , the relationship between $\mu$ and $\nu$ obeys $\mu=\frac{2}{\nu}$, by which we obtain $\nu=1.4674\pm 0.004$. Although the numerically obtained critical exponents $\mu$ and $\nu$ has slightly numerical deviation from the exact solution (with $\nu_{exact}=3/2$) of the single Rabi model HP ; ps2 , we argue that the quantum phase transition in the $\varphi=\frac{\pi}{2}$ is the same universality class as that in the single Rabi model. Within the neighborhood of the phase transition point, $\chi_{\mathcal{F}}$ scales as $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}=f\left[\eta^{1/\nu}\left(\mathcal{J}-\mathcal{J}_{max}\right)\right]$, where $f(x)$ is a universal scaling function. Under different $\eta$, we find the corresponding data of $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}$ are collapsed into one curve with the exact critical exponent $\nu=\frac{3}{2}$, signaling the same universal class as the single Rabi model HP ; ps2 . In fact, for the Dicke model Dicke_1 ; Dicke_2 and the Lipkin-Meshkov-Glick model LMG_1 ; LMG_2 ; LMG_3 , the critical exponent satisfies $3/2$ as well, therefore the conventional SPT of the non-Hermitian cascade Rabi cavities we have studied belongs to the same universality class as those in the above mentioned systems.
VI Experimental feasibility
The experimental exploration of the phenomena discussed here could be facilitated by a proposed circuit-QED architecture, as detailed in Ref. SC . This setup comprises two resonators, each equipped with an embedded artificial atom, specifically a superconducting flux qubit. The key feature of this design is the implementation of a coupling mechanism between the resonators that is both broadly and rapidly tunable. By leveraging fast-modulating fields, this tunable coupling can effectively emulate a phase-dependent interaction term, a crucial component for investigating PT-related physics. Furthermore, the architecture allows for the coupling between the flux qubits and the resonators to be pushed into the ultrastrong coupling regime ($g>0.1$), and even the deep-strong coupling regime ($g>1$). These advanced capabilities of the circuit-QED system provide a solid platform not only for testing the theoretical predictions presented in this article but also for realizing these concepts in practical, soon-to-be-available devices.
VII Summary
Herein, a non-Hermitian QRM consisting of two cascaded non-Hermitian Rabi cavities has been studied. Similar to Hermitian QRMs, the superradiance phase boundary of the non-Hermitian one can be exactly solved as well. Accompanied by the emergence of SPT, the ground state energy gets lower. Meanwhile, the atoms become active and radiating photons outward. However, the non-Hermiticity leads to the exotic phenomena without Hermitian counterpart. When the atom-optical field coupling phase satisfies $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, the superradiance phase boundary is surprisingly dominated by a constant independent of the strength of the atom-cavity field coupling, suggesting the unconventional SPT. For other coupling phases, the superradiance phase boundary depends on the cavity coupling strength and the atom-optica field coupling strength together, suggesting the conventional SPT. Meanwhile, we have argued that the conventional SPT belongs to the second-order quantum phase transition, while the unconventional SPT belongs to the first-order quantum phase transition. Even though the model we have studied is not $\mathcal{PT}$-symmetric, there still exists fully real energy spectrum. By the photon population, atom population, and the fidelity we have known that there is no quantum criticality for the unconventional SPT due to characteristics of the absence of asymptotic process and the single-parameter sharp drop of fidelity. Differently, for conventional SPT, there exists quantum criticality, and the critical exponent extracted from the scaling analyze of fidelity susceptibility show that the SPT belong to the same universality class as the single Rabi model. With the development of circuit-QED experimental technique, the found SPTs may be realized in high-performance circuit- QED architecture.
We acknowledge the support from NSFC under Grant No. 12174346.

### v2

@@
-Phase driven exact and unconventional superradiance phase transition in non-Hermitian cascaded Rabi cavities
+Phase driven unconventional superradiance phase transition in non-Hermitian cascaded quantum Rabi cavities
 Abstract
-This work reports the phase driven symmetry breaking and exact and unconventional superradiance phase transition in the non-Hermitian cascaded Rabi cavities.
-The non-Hermiticity is introduced in the coupling phase (denoted by $\varphi$) between the atom and the optical field.
-The exactness refers to the fact that the superradiance phase boundary is obtained analytically and verified by the observables.
-The unconventionality is reflected in that when $|\varphi|=\frac{\pi}{4}$ or $|\varphi|=\frac{3\pi}{4}$, the phase boundary is uniquely determined by $\mathcal{J}=\frac{1}{2}$ (where $\mathcal{J}$ is the dimensionless cavity coupling strength) and is independent of the atom-optical field coupling strength $g$.
-For other $\varphi$, the phase boundary is determined by $\mathcal{J}$ and the dimensionless atom-optical field coupling strength $g$ together.
-Besides, we find that there are phase driven first-order and second-order superradiance phase transitions, and the quantum criticality for the second-order superradiance phase transition is studied.
-In addition, the experimental feasibility is discussed.
-This work will stimulate the studies of non-Hermitian superradiance quantum phase transitions and their experimental realizations, as well as the underlying universality class of phase transitions.
+This study investigates phase-driven symmetry breaking leading to superradiance phase transitions in cascaded non-Hermitian quantum Rabi cavities.
+Non-Hermiticity is introduced via the phase coupling $\varphi$ between the atom and the optical field.
+In the thermodynamic limit of the quantum harmonic oscillator, we analytically derive the superradiance phase boundary, validated by observables.
+An unconventional quantum phase transition without a Hermitian analogue arises when $|\varphi|=\frac{\pi}{4}$ or $|\varphi|=\frac{3\pi}{4}$, where the phase boundary is uniquely determined by the cavity coupling, at $\mathcal{J}=\frac{1}{2}$, independent of the atom-photon coupling strength $g$.
+For other $\varphi$, the phase boundary relies on both $\mathcal{J}$ and $g$, similar to the scenario observed in Hermitian systems.
+Furthermore, we identify phase-driven first- and second-order superradiance phase transitions, focusing on the quantum criticality of the second-order transition by determining the critical exponents and the universality class.
+The feasibility of experimental realization is also discussed, aiming to inspire further studies on non-Hermitian superradiance quantum phase transitions.
 pacs:
@@
 I Introduction
-The quantum Rabi model (QRM) named as Rabi Rabi ; book describes the interaction of a two-level atom and a single-model optical field and can be viewed as the minimal version of the Dicke model Dicke .
-In the past years, there are efforts in finding the exactly analytically solutions of the QRM am_1 ; am_3 ; am_2 ; am_4 ; am_5 ; am_6 ; am_7 ; am_8 .
-Employing the Bargmann space representation bs , Braak obtained the transcendental function, namely the G-function, and found that the poles of the G-function were the exact solutions of QRM es1 .
-Quickly, Chen et al.
-obtained the exact solutions of a two-photon QRM by combining the Bogoliubov transformation and the G-function es2 .
-In the decade, the two methods have been used for calculating the exact solutions of the generalized QRMs ext_1 ; ext_2 ; ext_3 ; ext_4 ; ext_5 ; ext_6 ; ext_7 ; ext_8 ; ext_9 ; ext_10 ; ext_11 .
-It was found that the symmetry breaking will lead to the topological transition in the QRM tp1 and the generalized QRM tp2 .
-Accompanying by the search for the exact solutions, there are interests in studying the quantum criticality of the QRMs or the generalized QRMs HP ; ps1 ; ps2 ; R_d1 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 ; zhangyy_2 .
-Accompanied by the quantum phase transition, i.e., the superradiance phase transition (SPT), there will occur macroscopic photon excitations R_d1 .
-In 2015, Hwang et al.
-studied the universal dynamics and the critical behaviors of the QRM and the exact phase boundary of SPT was obtained by an effective low-energy approximation HP .
-They found that the critical exponent of SPT $\nu$ is $\nu=\frac{3}{2}$ HP .
-By analyzing a general scaling function, Liu et al.
-found that the $\nu$ in the anisotropic QRM sill satisfies $\nu=\frac{3}{2}$ and the phase boundary of SPT is obtained by a semi-class method ps2 .
-Quickly, Wei et al.
-numerically verified the exact $\nu$ by analyzing the critical behaviors of the finite frequency fidelity susceptibility qc_1 .
-Subsequent researches show that for a multi-cavity system with two qc_2 ; qc_3 or three zhangyy_2 cascaded QRMs, the crtical exponent is still equal to that of a single Rabi model, but for a three-level QRM, there occurs new critical exponent with $\nu=1$zhangyy_3 .
-We notice that the SPTs in the QRMs more or less depend on the atom-optical field coupling strength.
-It motivate us whether there exist an unconventional SPT which is independent of the atom-optical field coupling.
-Recently, the combination of the non-Hermiticity and the QRMs triggered intriguing physics.
-Cao et al.
-constructed a generalized non-Hermitian integrable QRM.
-They obtained the exact energy spectrum and eigenstates of this model by employing the Bethe ansatz es_NH_1 .
-The work provides a generalized strategy to engineer integrable non-Hermitian spin-boson models.
-By introducing the parity-time ($\mathcal{PT}$) symmetric gain and loss in the semi-classical Rabi model, Lee et al.
-PT_Rabi_1 found that there was $\mathcal{PT}$ symmetry breaking PT ; PT1 accompanied by the real-complex transition in energy.
-With the exact solutions of the $PT$-symmetric Rabi models, one can locate the $PT$-symmetry breaking point, i.e., the exceptional point (EP) EP .
-Xie et al.
-investigated the Floquet $\mathcal{PT}$-symmetric semi-classical Rabi model es_NH_2 .
-Under the multiple-photon resonance condition, they obtained the positions of the EPs from the exact Floquet spectrum.
-Employing the adiabatic approximation, Lu et al.
-obtained the exact energy spectrum of the $\mathcal{PT}$-symmetric QRM, by which the EPs were located es_NH_3 .
-With the emergence of EP, observables present discontinuous behaviors.
-By the Bogoliubov transformation and G-function, Li et al.
-solved the exact energy spectra for a one-photon and two-photon $\mathcal{PT}$-symmetric QRM, and the zeros of G-function located the EPs es_NH_4 .
-These works provide an inspiration for us to study whether there will be a completely real energy spectrum in a generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry.
-In addition, we will study the quantum criticality of the model and discuss the similarity and difference with Hermitian cases.
+The cornerstone model for light-matter interaction is the quantum Rabi model (QRM), which involves a two-level atom (qubit) coupled through dipole interaction to a quantized single-mode cavity field and can be seen as the minimal version of the Dicke modelRabi ; book ; Dicke .
+In recent years, there have been efforts to find exact analytical solutions for the QRM am_1 ; am_2 ; am_3 ; am_4 ; am_5 ; am_6 ; am_7 ; am_8 .
+Using the Bargmann space representation bs , Braak obtained the transcendental function, known as the G-function, and discovered that the poles of the G-function were the exact solutions of the QRM es1 .
+Shortly after, Chen et al.
+achieved the exact solutions of a two-photon QRM by combining the Bogoliubov transformation and the G-function es2 .
+Over the past decade, these methods have been employed to calculate the exact solutions of generalized QRMs ext_1 ; ext_2 ; ext_3 ; ext_4 ; ext_5 ; ext_6 ; ext_7 ; ext_8 ; ext_9 ; ext_10 ; ext_11 .
+It has been found that symmetry breaking can lead to topological transitions in both the QRM tp1 and generalized QRM tp2 .
+Parallel to the search for exact solutions, significant interest has emerged in exploring the quantum criticality of QRMs and generalized QRMs HP ; ps1 ; ps2 ; R_d1 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 ; zhangyy_2 .
+Quantum criticality denotes abrupt changes in a system’s properties at quantum phase transitions (QPT), where macroscopic observables display non-analytic behavior due to quantum fluctuations.
+In the context of light-matter interaction, this phenomenon is exemplified by quantum phase transitions such as the superradiance phase transition (SPT), characterized by macroscopic excitations in ground-state populationsreview .
+Experimental observations in ion traps experiment have confirmed a second-order QPT in the QRM, where researchers noted a rapid increase in both phonon number and internal ion excitations at the critical point.
+In the deep strong coupling regime of light-matter interaction models, a notable feature is the zero-temperature photon vacuum instability driving systems to undergo a QPT at critical coupling strength $g_{c}$.
+Below $g_{c}$, the system resides in a normal symmetric phase with the field in vacuum and atoms in their lowest electronic state.
+Above $g_{c}$, both the field and atoms become excited, indicating a superradiant phase with broken symmetry.
+The thermodynamic limit is approached by increasing the number of particles to infinity in the Dicke model, resembling a classical limit for the atomic part.
+In contrast, phase transitions in single-atom systems occur at the classical oscillator limit obtained by increasing the qubit/oscillator frequency ratio.
+Furthermore, the anisotropic quantum Rabi-Stark model (AQRSM), incorporating Stark coupling terms, demonstrates first-order and continuous QPTs xie2020first .
+Notably, excited-state quantum phase transitions (ESQPTs) have been extensively studied for both the standard quantum Rabi model (QRM) puebla2016excited and the AQRSM hu2023excited in the limit of infinite frequency ratio.
+These transitions involve non-analytic changes in the excited-state properties as a control parameter varies, occurring at finite energy densities and marked by singularities in the density of states.
+Studies have examined the universal dynamics and critical behaviors of the QRM, determining the exact phase boundary of the SPT using an effective low-energy approximation HP .
+The critical exponent of the SPT, $\nu$, was found to be $\nu=\frac{3}{2}$ HP .
+Analyzing a general scaling function has demonstrated that $\nu$ in the anisotropic QRM also satisfies $\nu=\frac{3}{2}$, and the phase boundary of the SPT was obtained via a semi-classical method ps2 .
+Subsequently, the exact $\nu$ was numerically confirmed by analyzing the critical behaviors of the finite frequency fidelity susceptibility qc_1 .
+Further research has revealed that for a multi-cavity system with two qc_2 ; qc_3 or three zhangyy_2 cascaded QRMs, the critical exponent remains the same as that of a single Rabi model.
+However, for a three-level QRM, a new critical exponent with $\nu=1$ emerges zhangyy_3 .
+Notably, SPTs in QRMs depend on the atom-photon coupling strength, motivating investigation into the existence of unconventional SPTs independent of this coupling.
+Recently, the combination of non-Hermiticity and QRMs has sparked intriguing developments.
+A generalized non-Hermitian integrable QRM was constructed, and its exact energy spectrum and eigenstates were obtained using the Bethe ansatz es_NH_1 .
+This work provides a generalized strategy to engineer integrable non-Hermitian spin-boson models.
+By introducing parity-time ($\mathcal{PT}$) symmetric gain and loss in the semi-classical Rabi model, it was discovered that $\mathcal{PT}$ symmetry breaking, accompanied by a real-complex transition in energy, occurs PT_Rabi_1 ; PT ; PT1 .
+With the exact solutions of the $\mathcal{PT}$-symmetric Rabi models, the $\mathcal{PT}$-symmetry breaking point, i.e., the exceptional point (EP), can be located EP .
+The Floquet $\mathcal{PT}$-symmetric semi-classical Rabi model was investigated, and under multiple-photon resonance conditions, the positions of the EPs from the exact Floquet spectrum were determined es_NH_2 .
+Using the adiabatic approximation, the exact energy spectrum of the $\mathcal{PT}$-symmetric QRM was obtained, locating the EPs es_NH_3 .
+With the emergence of EPs, observables present discontinuous behaviors.
+The exact energy spectra for one-photon and two-photon $\mathcal{PT}$-symmetric QRMs were solved using the Bogoliubov transformation and G-function, and the zeros of the G-function located the EPs es_NH_4 .
+Motivated by these studies, we will examine whether a fully real energy spectrum can emerge in a generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry.
+Furthermore, this paper explores the quantum criticality of the model, revealing unique behaviors and transitions that distinguish it from conventional systems.
+By comparing and contrasting these findings with the established features of Hermitian systems, we aim to illuminate the intriguing parallels and divergences between the two realms.
 The paper is organized as follows.
 In Sec.
-II we introduce the generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry.
+II, we introduce the generalized non-Hermitian QRM without $\mathcal{PT}$ symmetry.
 In Sec.
@@
 V, we discuss the quantum criticality of the SPT.
-A experimental feasibility is discussed in Sec.
+Experimental feasibility is discussed in Sec.
 VI, and a summary is presented in Sec.
@@
 $H=\sum_{j=1,2}h_{j}+J(a^{{\dagger}}_{1}+a_{1})(a^{{\dagger}}_{2}+a_{2}),$ (1)
-where $h_{j}$ describes non-Hermitian Rabi cavity $h_{j}=\omega_{p}a^{{\dagger}}_{j}a_{j}+\frac{\omega_{a}}{2}\sigma^{j}_{z}-\lambda e^{i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}$ with $j$ the cavity index and $a^{{\dagger}}_{j}$ ($a_{j}$) the creation (annihilation) of photons.
-$\omega_{p}$ is the optical field frequency, $\omega_{a}$ is the atomic frequency, and $\lambda e^{i\varphi}$ is the coupling strength between the atom and optical field with $\lambda$ the weight of strength and $\varphi$ the coupling phase.
-The non-Hermiticity of the system requires $\varphi\in(-\pi,0)\cup(0,\pi)$.
+where $h_{j}$ describes the non-Hermitian Rabi cavity
+$h_{j}=\omega_{p}a^{{\dagger}}_{j}a_{j}+\frac{\omega_{a}}{2}\sigma^{j}_{z}-\lambda e^{i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x},$
+with $j$ denoting the cavity index and $a^{{\dagger}}_{j}$ ($a_{j}$) representing the creation (annihilation) operators of photons.
+Here, $\omega_{p}$ is the optical field frequency, $\omega_{a}$ is the atomic frequency, and $\lambda e^{i\varphi}$ is the coupling strength between the atom and optical field, where $\lambda$ represents the strength and $\varphi$ is the coupling phase.
+The non-Hermiticity of the system imposes that $\varphi\in(-\pi,0)\cup(0,\pi)$.
 Figure 1: (Color online) Ground-state phase diagram in the $\mathcal{J}$-$g$ parameter space.
@@
 The blue solid lines denote the phase boundary, satisfying $1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)=2\mathcal{J}$.
-For a single non-Hermitian Rabi cavity, we argue that there is no parity-time ($\mathcal{PT}$) symmetry unless the coupling phase $\varphi=\pm\frac{\pi}{2}$.
-The time-reversal operator $\mathcal{T}$ dominates the complex conjugate operation, resulting in $\mathcal{T}a^{{\dagger}}_{j}(a_{j})\mathcal{T}$=$a^{{\dagger}}_{j}(a_{j})$.
-The parity operator $\mathcal{P}$ dominates the inversion operation, i.e., $\mathcal{P}\hat{x}_{j}\mathcal{P}$= $-\hat{x}_{j}$, and $\mathcal{P}\hat{p}_{j}\mathcal{P}$= $-\hat{p}_{j}$ PT .
-Thus, we have
+In the context of a single non-Hermitian Rabi cavity, parity-time ($\mathcal{PT}$) symmetry is absent unless the coupling phase attains the specific values $\varphi=\pm\frac{\pi}{2}$.
+The time-reversal operator $\mathcal{T}$ enacts a complex conjugation operation, yielding $\mathcal{T}a^{{\dagger}}_{j}(a_{j})\mathcal{T}=a^{{\dagger}}_{j}(a_{j})$.
+Similarly, the parity operator $\mathcal{P}$ performs an inversion transformation defined by $\mathcal{P}\hat{x}_{j}\mathcal{P}=-\hat{x}_{j}$ and $\mathcal{P}\hat{p}_{j}\mathcal{P}=-\hat{p}_{j}$ PT .
+Consequently, we have
 $\mathcal{PT}h_{j}\mathcal{TP}$ $=\omega_{p}a^{{\dagger}}_{j}a_{j}-\mathcal{P}\left[\lambda e^{-i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}\right]\mathcal{P}+\frac{\omega_{a}}{2}\sigma^{j}_{z}$ (2)
 $=\omega_{p}a^{{\dagger}}_{j}a_{j}+\lambda e^{-i\varphi}\left(a_{j}+a^{{\dagger}}_{j}\right)\sigma^{j}_{x}+\frac{\omega_{a}}{2}\sigma^{j}_{z}.$
-Only if $\varphi=\pm\frac{\pi}{2}$, the single non-Hermitian Rabi cavity is $\mathcal{PT}$-symmetric (see another prove in Ref. es_NH_4 ). For other values, it is not $\mathcal{PT}$-symmetric.
-For the cascaded non-Hermitian QRM, we demonstrate that this system is completely not $\mathcal{PT}$-symmetric.
-$\mathcal{P}$ dominates a mirror operation between the two cavities, i.e., $\mathcal{P}a^{{\dagger}}_{j}(a_{j})\mathcal{P}=a^{{\dagger}}_{j^{\prime}}(a_{j^{\prime}})$, $\mathcal{P}\sigma^{j}_{z}(\sigma^{j}_{x})\mathcal{P}=\sigma^{j^{\prime}}_{z}(\sigma^{j^{\prime}}_{x})$ (here $j\neq j^{\prime}$), and $\mathcal{T}$ leads to $\mathcal{T}i\mathcal{T}=-i$.
+Thus, the single non-Hermitian Rabi cavity manifests $\mathcal{PT}$-symmetry exclusively for $\varphi=\pm\frac{\pi}{2}$ (refer to an alternate proof in Ref.
+es_NH_4 ).
+For any other values of $\varphi$, $\mathcal{PT}$-symmetry is not exhibited.
+For the cascaded non-Hermitian QRM, we verify that this system is entirely not $\mathcal{PT}$-symmetric.
+The parity operator $\mathcal{P}$ mirrors operations between the two cavities, specifically $\mathcal{P}a^{{\dagger}}_{j}(a_{j})\mathcal{P}=a^{{\dagger}}_{j^{\prime}}(a_{j^{\prime}})$ and $\mathcal{P}\sigma^{j}_{z}(\sigma^{j}_{x})\mathcal{P}=\sigma^{j^{\prime}}_{z}(\sigma^{j^{\prime}}_{x})$ (where $j\neq j^{\prime}$), while the time-reversal operator $\mathcal{T}$ satisfies $\mathcal{T}i\mathcal{T}=-i$.
 Performing $\mathcal{PT}$ on the cavity coupling term, we have
 $\mathcal{PT}J\left(a^{{\dagger}}_{1}+a_{1}\right)\left(a^{{\dagger}}_{2}+a_{2}\right)\mathcal{TP}=J\left(a^{{\dagger}}_{1}+a_{1}\right)\left(a^{{\dagger}}_{2}+a_{2}\right).$ (3)
-Performing $\mathcal{PT}$ on $h_{j}$, we have
+Performing $\mathcal{PT}$ on $h_{j}$, we find that
 $\mathcal{PT}h_{j}\mathcal{TP}$ $=\omega_{p}a^{{\dagger}}_{j^{\prime}}a_{j^{\prime}}-\lambda e^{-i\varphi}\left(a_{j^{\prime}}+a^{{\dagger}}_{j^{\prime}}\right)\sigma^{j^{\prime}}_{x}+\frac{\omega_{a}}{2}\sigma^{j^{\prime}}_{z}$ (4)
-$\neq h_{j^{\prime}}.$
-Therefore, here the non-Hermitian cascaded QRM is not $\mathcal{PT}$-symmetric.
-III Exact and unconventional superradiance phase transition
+$\neq h_{j^{\prime}},$
+where $h_{j}$ is the Hamiltonian of the $j$-th cavity. Therefore, the cascaded non-Hermitian QRM does not exhibit $\mathcal{PT}$-symmetry.
+III Analytical Results
 To analytically solve the phase boundary of the ground-state superradiance phase transition, we introduce the dimensionless parameters $\eta=\frac{\omega_{a}}{\omega_{p}}$, $g=\frac{2\lambda}{\sqrt{\omega_{a}\omega_{p}}}$, and $\mathcal{J}=\frac{J}{\omega_{p}}$.
@@
 $\mathcal{H}=\frac{H}{\omega_{p}\eta}\equiv\frac{1}{2}\sum_{j=1,2}\left(x^{2}_{j}+\frac{y^{2}_{j}}{\eta^{2}}+\sigma^{j}_{z}-\sqrt{2}g\sigma^{j}_{x}x_{j}\right)+2\mathcal{J}x_{1}x_{2}.$ (8)
-As studied in Refs.
-ps1 ; ps2 , the SPT appears in the under the infinite $\eta$ limit.
-Under this condition, we can obtain the analytical phase boundary of SPT.
-When $\eta\rightarrow\infty$, $\mathcal{H}$ reduces to
-$\mathcal{H}=\frac{1}{2}\sum_{j=1,2}\left[x^{2}_{j}+\sigma^{j}_{z}-\sqrt{2}g\sigma^{j}_{x}x_{j}\right]+2\mathcal{J}x_{1}x_{2}.$ (9)
-By diagonalizing $\mathcal{H}$, we obtain the energies $E_{\pm}(x_{1},x_{2})$ as
-$E_{\pm}(x_{1},x_{2})=\frac{1}{2}\sum_{j=1,2}\left(x^{2}_{j}\pm\sqrt{1+2g^{2}x^{2}_{j}}\right)+2\mathcal{J}x_{1}x_{2}.$ (10)
-Performing the Taylor expansion on $E_{\pm}(x_{1},x_{2})$, and truncating them to the quadratic term $\mathcal{O}(x^{4}_{j})$, the low-energy branch $E_{-}(x_{1},x_{2})$ (contains the ground-state information) becomes
+As studied in Refs. ps1 ; ps2 , the superradiance phase transition (SPT) emerges under the infinite $\eta$ limit. Under this condition, we derive the analytical form of the Hamiltonian $\mathcal{H}$:
+$\mathcal{H}=\frac{1}{2}\sum_{j=1,2}\left[x_{j}^{2}+\sigma_{z}^{j}-\sqrt{2}g\sigma_{x}^{j}x_{j}\right]+2\mathcal{J}x_{1}x_{2}.$ (9)
+Diagonalizing $\mathcal{H}$, we obtain the energies $E_{\pm}(x_{1},x_{2})$:
+$E_{\pm}(x_{1},x_{2})=\frac{1}{2}\sum_{j=1,2}\left(x_{j}^{2}\pm\sqrt{1+2g^{2}x_{j}^{2}}\right)+2\mathcal{J}x_{1}x_{2}.$ (10)
+Expanding $E_{\pm}(x_{1},x_{2})$ to quadratic terms $\mathcal{O}(x_{j}^{4})$, the low-energy branch $E_{-}(x_{1},x_{2})$ (representing the ground-state information) simplifies to:
 $E_{-}(x_{1},x_{2})=\frac{1}{2}\psi^{T}\Lambda\psi,$ (11)
-in which $\psi=(x_{1},x_{2})^{T}$ and
+where $\psi=(x_{1},x_{2})^{T}$ and $\Lambda$ is defined as:
 $\Lambda=\begin{pmatrix}1-g^{2}e^{i2\varphi}&2\mathcal{J}\\ 2\mathcal{J}&1-g^{2}e^{i2\varphi}\end{pmatrix}.$ (12)
-Evidently, the matrix $\Lambda$ consists of the real part ${\rm Re}(\Lambda)$ and the pure imaginary part ${\rm Im}(\Lambda)$ (which can be considered as the background loss), i.e., $\Lambda={\rm Re}(\Lambda)+{\rm Im}(\Lambda)$, with
+The matrix $\Lambda$ separates into its real part ${\rm Re}(\Lambda)$ and pure imaginary part ${\rm Im}(\Lambda)$, representing the background loss:
 ${\rm Re}(\Lambda)$ $=\left[1-g^{2}(\cos^{2}\varphi-\sin^{2}\varphi)\right]\sigma_{0}+2\mathcal{J}\sigma_{x},$ (13)
@@
 The photonic truncation number is $N=80$.
-In fact, the phase boundary of SPT is nothing to do with background loss ${\rm Im}(\Lambda)$, but is determined by ${\rm Re}(\Lambda)$.
-After diagonalizing ${\rm Re}(\Lambda)$, the eigenvalues ϵ±subscriptitalic-$\epsilon_{\pm}$ of ${\rm Re}(\Lambda)$ is obtained as $\epsilon_{\pm}=1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)\pm 2\mathcal{J}$.
-When the low-energy branch in ϵ±subscriptitalic-$\epsilon_{\pm}$ equals to zero, i.e, $\epsilon_{-}=0$, we can extract the phase boundary of the superradiance phase transition, and it satisfies $1-g^{2}\left(\cos^{2}\varphi-\sin^{2}\varphi\right)=2\mathcal{J}$.
-From the expression of the phase boundary, we can see that there are special phases, i.e., $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, at which the phase boundary is fixed at $\mathcal{J}=\frac{1}{2}$, and is independent of the atom-optical field coupling parameter $g$.
-Such type of SPT independent of atom-optical field coupling is counter-intuitive and has not been reported before, so we name this type of SPT as unconventional SPT.
-While $\gamma$ is taken at other values, the phase boundary depends on the $g$ and $\mathcal{J}$ together.
-Thus, we name this atom-optical field coupling dependent SPT as the conventional SPT.
-Without loss of generality, in the following, we will study the energy, photon population, qubit population, the fidelity and the fidelity susceptibility in the ground state at $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ to characterize the properties of the SPT of the non-Hermitian two-Rabi cavity system.
-The ground state here refers to the state with the smallest real part of the energy.
-Intuitively, the phase diagrams for $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ are plotted in Fig.
-1(a) and 1(b), respectively.
-The red regions denote the superradiance phases, separated from the normal phases (yellow regions) by the phase boundaries (blues curves).
+The phase boundary of the superradiance phase transition (SPT) is primarily determined by ${\rm Re}(\Lambda)$ rather than ${\rm Im}(\Lambda)$.
+Diagonalizing ${\rm Re}(\Lambda)$ yields the eigenvalues $\epsilon_{\pm}=1-g^{2}(\cos^{2}\varphi-\sin^{2}\varphi)\pm 2\mathcal{J}$.
+The condition for the SPT phase boundary, where the low-energy branch $\epsilon_{-}=0$, results in $1-g^{2}(\cos^{2}\varphi-\sin^{2}\varphi)=2\mathcal{J}$.
+This phase boundary reveals special phases at $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, where $\mathcal{J}=\frac{1}{2}$ independently of the atom-photon coupling parameter $g$.
+Such a phase transition, independent of the coupling $g$, is termed unconventional SPT as it has not been reported previously.
+For other values of $\varphi$, the phase boundary depends jointly on $g$ and $\mathcal{J}$, categorizing it as conventional SPT.
+The subsequent analysis examines ground state energies, photon and qubit populations, fidelity, and fidelity susceptibility at $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$ to characterize the properties of the non-Hermitian two-Rabi cavity system.
+Here, the ground state refers to the state with the lowest real part of the energy.
+Figures 1(a) and 1(b) illustrate phase diagrams for $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$, respectively, where red regions indicate superradiance phases separated by blue curves representing phase boundaries, and normal phases are depicted in yellow regions.
 Figure 3: (Color online) The ground-state photon population $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ as a function of $\mathcal{J}$.
@@
 IV Signatures of the superradiance phase transition
-We detect the signatures and check the phase boundary of the SPT.
-We start by studying the ground-state energy to reveal the SPT.
-The energies are sorted by the real parts of the energies, and the ground-state energy means the energy with the lowest real part.
-Under $\varphi=\frac{\pi}{2}$ and different $\eta$, the real parts of the ground-state energies ${\rm Re}(E_{0})/\eta$ as a function of $\mathcal{J}$ for $g=0.7$ are plotted in Fig.
+To identify the signatures and establish the phase boundary of the Superradiance Phase Transition (SPT), we commence our investigation by analyzing the ground-state energy.
+Energies are ranked based on their real parts, with the ground-state energy denoting the lowest real part.
+Under the condition of $\varphi=\frac{\pi}{2}$ and varying $\eta$, the real parts of the ground-state energies $\mathrm{Re}(E_{0})/\eta$ as a function of $\mathcal{J}$ for $g=0.7$ are depicted in Fig.
 2(a).
-The photonic truncation number is taken at $N=80$.
-As can be seen that with the emergence of SPT, ${\rm Re}(E_{0})/\eta$ for different $\eta$ decrease from $-1$ to lower values.
-According to the obtained analytical result, the transition point under $\varphi=\frac{\pi}{2}$ and $g=0.7$ should be at $\mathcal{J}=0.745$.
-Intuitively, the SPT occurs at the transition points predicted by the exactly analytical result (shown with the vertical black line).
-Meanwhile, we calculate ${\rm Re}(E_{0})/\eta$ in the $\varphi=\frac{\pi}{4}$ case and ${\rm Re}(E_{0})/\eta$-$\mathcal{J}$ relations for $g=0.4$ are plotted in Fig.
-2(b).
-Similar to the $\varphi=\frac{\pi}{2}$, when $\mathcal{J}$ exceeds to the transition point, ${\rm Re}(E_{0})/\eta$ for different $\eta$ in the $\varphi=\frac{\pi}{4}$ case decrease from $-1$ to lower values, signaling the appearance of SPT as well.
-From the analytical expression of phase boundary, it is known that the transition point under $\varphi=\frac{\pi}{2}$ should be at $J=0.5$ regardless of what value of $g$ is.
-It is readily seen that for $g=0.4$, the transition points of ${\rm Re}(E_{0})/\eta$ occur at the analytically obtained transition points.
-Although the ground state energy of the system decreases after the occurrence of conventional and unconventional SPTs, it should be pointed out that they do not belong to the same phase transition type.
-As shown in Fig.
-1(c), the second-order derivative of ${\rm Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{2}$ and $g=0.7$ is not continuous at the the transition point.
-Figure 1(d) shows that the first-order derivative of ${\rm Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{4}$ and $g=0.4$.
-The results show that the conventional SPT should be the second-order quantum phase transition, while the unconventional SPT should be the first-order quantum phase transition.
-In fact, only the $|\varphi|=\frac{\pi}{4}$ and $|\varphi|=\frac{3\pi}{4}$ cases correspond to the second-order quantum phase transition, while the cases with other $\varphi$ correspond to the first-order quantum phase transition.
-However, we are more concerned with conventional and unconventional superradiance phase transitions in this paper, so the results of other $\varphi$ are not considered
+Here, the photonic truncation number $N=80$ ensures the convergence of results.
+It is observed that with the onset of the SPT, $\mathrm{Re}(E_{0})/\eta$ for different $\eta$ decreases from $-1$ to lower values.
+According to the analytical predictions, the transition point under $\varphi=\frac{\pi}{2}$ and $g=0.7$ is expected to occur at $\mathcal{J}=0.745$, marked by the vertical black line.
+This confirms the occurrence of the SPT at the theoretically derived transition points.
+Similarly, the plot in Fig.
+2(b) illustrates the relationship between $\mathrm{Re}(E_{0})/\eta$ and $\mathcal{J}$ for $\varphi=\frac{\pi}{4}$ with $g=0.4$.
+As in the $\varphi=\frac{\pi}{2}$ case, when $\mathcal{J}$ surpasses the transition point, $\mathrm{Re}(E_{0})/\eta$ for different $\eta$ decreases from $-1$ to lower values, indicating the onset of the SPT.
+The analytical phase boundary indicates that the transition point under $\varphi=\frac{\pi}{4}$ should be at $\mathcal{J}=0.5$, which aligns with the observed transitions for $g=0.4$.
+The analysis of Fig.
+1(c) shows that the second-order derivative of $\mathrm{Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{2}$ and $g=0.7$ is discontinuous at the transition point, characteristic of a second-order quantum phase transition.
+Conversely, Fig.
+1(d) demonstrates that the first-order derivative of $\mathrm{Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{4}$ and $g=0.4$.
+These findings indicate that the conventional SPT is a second-order quantum phase transition, while the unconventional SPT is a first-order quantum phase transition.
+Specifically, cases where $|\varphi|=\frac{\pi}{4}$ and $|\varphi|=\frac{3\pi}{4}$ correspond to second-order quantum phase transitions, whereas other values of $\varphi$ correspond to first-order quantum phase transitions.
+However, this study focuses primarily on conventional and unconventional superradiance phase transitions, thus excluding detailed consideration of results for other $\varphi$ values.
 Figure 4: (Color online) The ground-state atom population $\sigma^{1(2)}_{z}$ as a function of $\mathcal{J}$.
@@
 We can see that there are macroscopic photon excitations when $\mathcal{J}$ is larger than $0.5$, while the photon excitations are evidently suppressed when $J<0.5$.
-A reasonable explaination for the macroscopic photon excitations in the suprradiance phase is that the atoms are excited and become active.
-To check this point, the atom populations $\langle\sigma^{1(2)}_{z}\rangle$ as a function of $\mathcal{J}$ for different $\varphi$ and $g$ are plotted in Fig.
+In the superradiance phase, in addition to the macroscopic excitations of the photons, the atoms also become excited and active.
+To verify this, we plot the populations of atoms $\langle\sigma^{1(2)}_{z}\rangle$ as a function of $\mathcal{J}$ for different $\varphi$ and $g$ in Fig.
 4(a)-4(d).
-We can see that when $\mathcal{J}$ is smaller than the transition point, the atoms are stable in the ground state, with populations $\langle\sigma^{1(2)}_{z}\rangle=-1$.
-When $\mathcal{J}$ crosses the transition point, it is seen that the populations of atoms deviate from -1, implying that the atoms are excited and become active.
-The excited atoms dominate the macroscopic photon excitations in the SPT.
-Although the photon excitations and the atom populations of the system vary obviously with occurrence of SPT, the two types of SPTs lead to different varying behaviors of observables.
-From Fig.
-3, we can see that when the conventional SPT happens, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ continuously varies to macroscopic value, while the unconventional SPT happens, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ directly jumps to macroscopic values without asymptotic process.
-Similarly, we can see the obvious difference between the two types of SPTs from the atomic populations in Fig.
+When $\mathcal{J}$ is below the transition point, the atoms remain in their ground state with populations $\langle\sigma^{1(2)}_{z}\rangle=-1$.
+However, as $\mathcal{J}$ crosses the transition point, the populations deviate from -1, indicating that the atoms are excited and active.
+These excited atoms drive the macroscopic photon excitations observed in the SPT.
+While the photon excitations and atom populations of the system change noticeably with the occurrence of SPT, the two types of SPTs exhibit distinct observable behaviors.
+As illustrated in Fig.
+3, during a conventional SPT, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ continuously varies to macroscopic values.
+In contrast, during an unconventional SPT, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ abruptly jumps to macroscopic values without an asymptotic process.
+This distinction is similarly evident in the atomic populations shown in Fig.
 4.
@@
 The photonic truncation number is $N=36$.
-In addition to the photon population and the atom population, the excitation gaps behind the two types of SPTs are different as well.
-The excitation gaps here contain two meanings.
-One is the real excitation gap ${\rm Re}(\Delta_{E})/\eta$ between the real parts of the lowest two energies.
-Another is the imaginary excitation gap ${\rm Im}(\Delta_{E})/\eta$ between the imaginary parts of the lowest two energies.
-Figures.
-5(a1)-5(d1) (Figs.
-5(a2)-5(d2)) presents ${\rm Re}(\Delta_{E})/\eta$ (${\rm Im}(\Delta_{E})/\eta$) as a function of $\mathcal{J}$ under different $\eta$.
-We can see that no matter $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$, the real excitation gap experiences a gap closing process when $\mathcal{J}$ exceeds the transition point.
-However, for the imaginary excitation gap, there is an obvious difference between the two studied cases.
-From Figs.
-5(a2) and 5(b2), we can see that ${\rm Im}(\Delta_{E})/\eta$ always keeps zero (${\rm Im}(\Delta_{E})/\eta$ far less than $10^{-14}$ can be numerically regarded as zero.) before and after the transition point.
-From Figs.
-5(c2) and 5(d2), we can see that the imaginary excitation gaps experience the gap closing process, similar to the circumstances for the real excitation gaps.
+In addition to the photon and atom populations, the excitation gaps associated with the two types of SPTs also exhibit distinct behaviors.
+These excitation gaps have two components: the real excitation gap ${\rm Re}(\Delta_{E})/\eta$, which is the gap between the real parts of the lowest two energies, and the imaginary excitation gap ${\rm Im}(\Delta_{E})/\eta$, which is the gap between the imaginary parts of the lowest two energies.
+Figures 5(a1)-5(d1) show ${\rm Re}(\Delta_{E})/\eta$ as a function of $\mathcal{J}$ for different $\eta$, while Figures 5(a2)-5(d2) present ${\rm Im}(\Delta_{E})/\eta$ under the same conditions.
+Regardless of whether $\varphi=\frac{\pi}{2}$ or $\varphi=\frac{\pi}{4}$, the real excitation gap undergoes a closing process when $\mathcal{J}$ surpasses the transition point.
+However, a noticeable difference arises in the imaginary excitation gaps between the two cases.
+As seen in Figures 5(a2) and 5(b2), ${\rm Im}(\Delta_{E})/\eta$ remains zero (values much smaller than $10^{-14}$ can be numerically considered as zero) both before and after the transition point.
+In contrast, Figures 5(c2) and 5(d2) demonstrate that the imaginary excitation gaps experience a closing process similar to that of the real excitation gaps.
 Figure 6: (Color online) The maximal imaginary part of the energy spectrum ${\rm Im}(E)_{max}/\eta$ as functions of $\mathcal{J}$ under different $\eta$.
@@
 The photonic truncation number is $N=36$.
-It is noted that for a $\mathcal{PT}$-symmetric Rabi system, there exists real-complex transition in energy spectrum PT ; PT1 ; es_NH_2 ; es_NH_3 ; es_NH_4 .
-Surprisingly, for the non-Hermitian two-Rabi cavity system without $\mathcal{PT}$ symmetry, there exists real energy spectrum as well.
-Under $\varphi=\frac{\pi}{2}$, the maximal imaginary of the energy spectra for $g=0.7$ and $g=0.8$, i.e., ${\rm Im}(E)_{max}$ are plotted in Figs.
+For a $\mathcal{PT}$-symmetric Rabi system, a real-to-complex transition in the energy spectrum is well-established PT ; PT1 ; es_NH_2 ; es_NH_3 ; es_NH_4 .
+Remarkably, the non-Hermitian two-Rabi cavity system without $\mathcal{PT}$ symmetry also exhibits a real energy spectrum.
+Under the condition $\varphi=\frac{\pi}{2}$, the maximum imaginary part of the energy spectra for $g=0.7$ and $g=0.8$, denoted as ${\rm Im}(E)_{\text{max}}$, is plotted in Figs.
 6(a) and 6(b), respectively.
-The analytically obtained superradiance transition point for $g=0.7$ is $J=0.745$, and that for $g=0.8$ is $J=0.82$.
-${\rm Im}(E)_{max}/\eta$ approaches to zero before and after the superradiance transition point implying that the energy spectra under $\varphi=\frac{\pi}{2}$ are real, and there is no real-complex transition in energy spectrum.
-Compared to the $\varphi=\frac{\pi}{2}$ case, the spectral property of the $\varphi=\frac{\pi}{4}$ is different.
-Under $\varphi=\frac{\pi}{4}$, ${\rm Im}(E)_{max}/\eta$ for $g=0.4$ and $g=0.6$ are plotted in Figs.
-6(c) and 6(d), respectively.
-Different from the $\varphi=\frac{\pi}{2}$ cases, here the energy spectra are complex because of the nonvanishing ${\rm Im}(E)_{max}/\eta$.
-In fact, the real and complex characteristics of the energy spectrum can be understood analytically as well.
-From Eq.
+The analytically determined superradiance transition point is $\mathcal{J}=0.745$ for $g=0.7$ and $\mathcal{J}=0.82$ for $g=0.8$.
+${\rm Im}(E)_{\text{max}}/\eta$ approaches zero both before and after the superradiance transition point, indicating that the energy spectra are real and no real-to-complex transition occurs.
+In contrast, the spectral properties under $\varphi=\frac{\pi}{4}$ are distinct.
+Figures 6(c) and 6(d) show ${\rm Im}(E)_{\text{max}}/\eta$ for $g=0.4$ and $g=0.6$, respectively.
+Unlike the $\varphi=\frac{\pi}{2}$ case, the energy spectra under $\varphi=\frac{\pi}{4}$ are complex due to the non-zero ${\rm Im}(E)_{\text{max}}/\eta$.
+This distinction can be analytically understood from Eq.
 (10) and Eq.
-(13), we know that when $\varphi=\frac{\pi}{2}$, the imaginary part disappears, but when $\varphi=\frac{\pi}{4}$, the imaginary part still exists.
-The results in Fig.
-6 provide numerical validation of the previous analytical derivation.
-Considering the fact that the $\varphi=\frac{\pi}{2}$ cases have full real energy spectra, we conjecture that here the conventional SPT belongs to the same universality class as the Hermitian Rabi systems HP ; ps2 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 and the critical exponent of the phase transition can be obtained from the critical behavior.
-However, for the complex energy spectrum case ($\varphi=\frac{\pi}{4}$), we conjecture that there is no critical behavior.
-Although the existence of quantum criticality of conventional SPT and the absence of quantum criticality of unconventional SPT have been reflected in the photon population and atom population, we still want to further verify them by studying the ground-state fidelity and fidelity susceptibility.
+(13).
+When $\varphi=\frac{\pi}{2}$, the imaginary part vanishes, whereas for $\varphi=\frac{\pi}{4}$, it persists.
+The numerical results in Fig.
+6 confirm these analytical insights.
+The full real nature of the energy spectra for $\varphi=\frac{\pi}{2}$ suggests that the conventional SPT in this case shares the universality class with Hermitian QRM systems HP ; ps2 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 , allowing for derivation of the critical exponent from its critical behavior.
+Conversely, for the complex energy spectrum case ($\varphi=\frac{\pi}{4}$), we hypothesize the absence of critical behavior.
+Although quantum criticality has been reflected in photon and atom populations for conventional SPT and its absence for unconventional SPT, we aim to validate these conjectures further through examination of ground-state fidelity and fidelity susceptibility.
 Figure 7: (Color online) Fidelity susceptibility $\chi_{\mathcal{F}}$ as functions of $\mathcal{J}$ under different $\eta$ with $\varphi=\frac{\pi}{2}$ and $g=0.7$ in (a), and with $\varphi=\frac{\pi}{2}$ and $g=0.8$ in (b).
@@
 The photonic truncation number in all calculations is $N=80$.
-V conventional and unconventional superradiance phase transition
-We focus on the quantum phase transition and its critical behavior triggered by the cavity coupling parameter $\mathcal{J}$.
-The fidelity for the ground state $\mathcal{F}$ is defined as fidelity_0 ; fidelity_1 ; fidelity_2 ; fidelity_3
-$\mathcal{F}=|\langle\psi_{0}(\mathcal{J})|\psi_{0}(\mathcal{J}+\delta\mathcal{J})|,$ (14)
-which weighs the survival probability of the ground state after the cavity coupling parameter $\mathcal{J}$ from an infinitesimal distance $\delta\mathcal{J}$.
-Further, according to Refs.
-FS_1 ; FS_2 ; FS_3 ; FS_4 ; FS_5 , we can obtain the fidelity susceptibility $\chi_{\mathcal{F}}$ as
+V Comparative Study of Conventional vs. Unconventional Superradiance Phase Transitions
+We focus on the quantum phase transition induced by the cavity coupling parameter $\mathcal{J}$.
+The fidelity for the ground state $\mathcal{F}$ is defined as fidelity_0 ; fidelity_1 ; fidelity_2 ; fidelity_3 :
+$\mathcal{F}=|\langle\psi_{0}(\mathcal{J})|\psi_{0}(\mathcal{J}+\delta\mathcal{J})|\rangle|,$ (14)
+which measures the overlap of the ground state before and after a small change $\delta\mathcal{J}$ in $\mathcal{J}$.
+According to Refs.
+FS_1 ; FS_2 ; FS_3 ; FS_4 ; FS_5 , the fidelity susceptibility $\chi_{\mathcal{F}}$ is given by:
 $\chi_{\mathcal{F}}=\lim_{\delta\mathcal{J}\rightarrow 0}\frac{-2\ln\mathcal{F}(\mathcal{J},\delta\mathcal{J})}{(\delta\mathcal{J})^{2}}.$ (15)
-If there exists quantum phase transition, the fidelity will experience a drop at the phase transition point, and the fidelity susceptibility will peak at the phase transition point.
+If a quantum phase transition exists, $\mathcal{F}$ exhibits a drop at the transition point, while $\chi_{\mathcal{F}}$ peaks.
 Figure 8: (Color online) (a) Universal finite $\eta$ scaling relation of the maximal fidelity susceptibility $\chi^{max}_{F}$.
@@
 The involved parameters are $\varphi=\frac{\pi}{2},$$g=0.7$, and $\delta J=10^{-5}$.
-Under $\varphi=\frac{\pi}{2}$, $\chi_{\mathcal{F}}/\eta$ for $g=0.7$ and $g=0.8$ are plotted in Figs.
+Under $\varphi=\frac{\pi}{2}$, $\chi_{\mathcal{F}}/\eta$ for $g=0.7$ and $g=0.8$ are shown in Figs.
 7(a) and 7(b), respectively.
-As seen that with the increase of $\mathcal{J}$, $\chi_{\mathcal{F}}/\eta$ smoothly reaches the peak and then smoothly decreases.
-Besides, $\chi_{\mathcal{F}}/\eta$ presents obvious critical behaviors, because it is seen that with the increase of $\eta$, the peak of $\chi_{\mathcal{F}}$ at the coupling parameter $\mathcal{J}_{max}$ gradually tends to the exact solution.
-Moreover, with the increase of $\eta$, the peak of $\chi_{\mathcal{F}}$ get higher.
-However, for the $\varphi=\frac{\pi}{4}$ cases, there is no critical phenomenon.
-Under $\varphi=\frac{\pi}{4}$, the fidelity $\mathcal{F}$ for $g=0.4$ and $g=0.6$ are plotted in Figs.
-7(c) and 7(d), respectively.
-Intuitively, $\mathcal{F}$ sharply drop to zero at single parameter $\mathcal{J}$ and there is no asymptotic behavior between the valley of $\mathcal{F}$ (or the peak of $\chi_{\mathcal{F}}$) and $\eta$.
-In other words, the conventional SPT exists quantum criticality while there is no quantum criticality for the unconventional SPT.
-Next, we study the quantum criticality of the conventional SPT ($\varphi=\frac{\pi}{2}$ case).
-From the $\chi_{\mathcal{F}}$-$\eta$ curves, the maximal fidelity susceptibility $\chi^{max}_{F}$ under a specific $\eta$ is obtained.
-Then, the universal finite $\eta$ scaling relation of $\chi_{\mathcal{F}}$ for $g=0.7$ is plotted in Fig.
+As $\mathcal{J}$ increases, $\chi_{\mathcal{F}}/\eta$ smoothly reaches a peak and subsequently decreases.
+Notably, $\chi_{\mathcal{F}}/\eta$ displays clear critical behavior; as $\eta$ increases, the peak at the coupling parameter $\mathcal{J}_{max}$ approaches the exact solution.
+Conversely, for $\varphi=\frac{\pi}{4}$, no critical phenomena are observed.
+As seen in Figs.
+7(c) and 7(d), the fidelity curves sharply drop to zero at a single $\mathcal{J}$ value without any asymptotic behavior between the fidelity valley (or $\chi_{\mathcal{F}}$ peak) and $\eta$.
+This indicates the presence of quantum criticality in conventional SPT and its absence in unconventional SPT.
+Next, we analyze the quantum criticality of conventional SPT ($\varphi=\frac{\pi}{2}$).
+From the $\chi_{\mathcal{F}}$ vs.
+$\eta$ curves, we extract the maximal fidelity susceptibility $\chi^{max}_{F}$ for a specific $\eta$.
+The universal finite $\eta$ scaling relation of $\chi_{\mathcal{F}}$ for $g=0.7$ is depicted in Fig.
 8(a).
-The black fitting curve satisfies $\ln\chi_{\mathcal{F}}=(-1.363\pm 0.004)\ln\eta+2.388$.
-It implies the scaling exponent $\mu$ of $\chi_{\mathcal{F}}$ is $\mu=-1.363\pm 0.004$.
-As introduced in Ref.
-qc_1 ; RMP2011 , the relationship between $\mu$ and $\nu$ obeys $\mu=\frac{2}{\nu}$, by which we obtain $\nu=1.4674\pm 0.004$.
-Although the numerically obtained critical exponents $\mu$ and $\nu$ has slightly numerical deviation from the exact solution (with $\nu_{exact}=3/2$) of the single Rabi model HP ; ps2 , we argue that the quantum phase transition in the $\varphi=\frac{\pi}{2}$ is the same universality class as that in the single Rabi model.
-Within the neighborhood of the phase transition point, $\chi_{\mathcal{F}}$ scales as $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}=f\left[\eta^{1/\nu}\left(\mathcal{J}-\mathcal{J}_{max}\right)\right]$, where $f(x)$ is a universal scaling function.
-Under different $\eta$, we find the corresponding data of $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}$ are collapsed into one curve with the exact critical exponent $\nu=\frac{3}{2}$, signaling the same universal class as the single Rabi model HP ; ps2 .
-In fact, for the Dicke model Dicke_1 ; Dicke_2 and the Lipkin-Meshkov-Glick model LMG_1 ; LMG_2 ; LMG_3 , the critical exponent satisfies $3/2$ as well, therefore the conventional SPT of the non-Hermitian cascade Rabi cavities we have studied belongs to the same universality class as those in the above mentioned systems.
+The fitted black curve satisfies $\ln\chi_{\mathcal{F}}=(-1.363\pm 0.004)\ln\eta+2.388$, indicating a scaling exponent $\mu=-1.363\pm 0.004$.
+Following the relationship $\mu=\frac{2}{\nu}$ qc_1 ; RMP2011 , we derive $\nu=1.4674\pm 0.004$.
+Although numerically slightly deviating from the exact solution ($\nu_{exact}=3/2$) of the single Rabi model HP ; ps2 , we argue that the quantum phase transition under $\varphi=\frac{\pi}{2}$ shares the same universality class as the single Rabi model.
+Around the transition point, $\chi_{\mathcal{F}}$ scales as $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}=f\left[\eta^{1/\nu}\left(\mathcal{J}-\mathcal{J}_{max}\right)\right]$, where $f(x)$ is a universal scaling function.
+From Fig.
+8(b), across various $\eta$, we observe that the data collapse into a single curve with the critical exponent $\nu=\frac{3}{2}$, indicating the same universality class as the single Rabi model HP ; ps2 .
+Indeed, similar critical exponents ($\nu=3/2$) are observed in the Dicke model Dicke_1 ; Dicke_2 and the Lipkin-Meshkov-Glick model LMG_1 ; LMG_2 ; LMG_3 , further affirming that the conventional SPT in the non-Hermitian cascade Rabi cavities studied here belongs to this universal class.
 VI Experimental feasibility
@@
 These advanced capabilities of the circuit-QED system provide a solid platform not only for testing the theoretical predictions presented in this article but also for realizing these concepts in practical, soon-to-be-available devices.
+Similarly, experiments exploring our findings can also be conducted in ion trap setups.
+Ion traps utilize individual ions confined in harmonic traps as qubits, offering precise control over quantum states and interactions.
+This setup enables the emulation of phase-dependent interactions required to investigate PT-related physics, similar to circuit-QED architectures.
+The trapped-ion system offers a versatile platform combining laser-coupled spin and bosonic degrees of freedom, ideal for simulating Hamiltonians of light-matter interactions.
+Demonstrated capabilities include quantum simulations of many-body spin models experiment2 , single-ion quantum Rabi models experiment , and recently, the Rabi-Hubbard (RH) model with up to 16 ions experiment3 , exploring equilibrium phase transitions and quantum dynamical properties using spin observables.
+The scalability and controllability of ion trap systems make them suitable for exploring both conventional and unconventional superradiance phase transitions in non-Hermitian quantum Rabi-dimer models and other related phenomena.
 VII Summary
-Herein, a non-Hermitian QRM consisting of two cascaded non-Hermitian Rabi cavities has been studied.
-Similar to Hermitian QRMs, the superradiance phase boundary of the non-Hermitian one can be exactly solved as well.
-Accompanied by the emergence of SPT, the ground state energy gets lower.
-Meanwhile, the atoms become active and radiating photons outward.
-However, the non-Hermiticity leads to the exotic phenomena without Hermitian counterpart.
-When the atom-optical field coupling phase satisfies $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, the superradiance phase boundary is surprisingly dominated by a constant independent of the strength of the atom-cavity field coupling, suggesting the unconventional SPT.
-For other coupling phases, the superradiance phase boundary depends on the cavity coupling strength and the atom-optica field coupling strength together, suggesting the conventional SPT.
-Meanwhile, we have argued that the conventional SPT belongs to the second-order quantum phase transition, while the unconventional SPT belongs to the first-order quantum phase transition.
-Even though the model we have studied is not $\mathcal{PT}$-symmetric, there still exists fully real energy spectrum.
-By the photon population, atom population, and the fidelity we have known that there is no quantum criticality for the unconventional SPT due to characteristics of the absence of asymptotic process and the single-parameter sharp drop of fidelity.
-Differently, for conventional SPT, there exists quantum criticality, and the critical exponent extracted from the scaling analyze of fidelity susceptibility show that the SPT belong to the same universality class as the single Rabi model.
-With the development of circuit-QED experimental technique, the found SPTs may be realized in high-performance circuit- QED architecture.
+Exploring a non-Hermitian Quantum Rabi Model composed of two cascaded non-Hermitian Rabi cavities, we investigate its superradiance phase boundary akin to its Hermitian counterpart.
+As the system undergoes SPT, transitioning from a quantum vacuum state to one that becomes macroscopically populated, we observe unique exotic phenomena arising from non-Hermitian effects.
+Notably, for specific atom-photon coupling phases such as $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, the superradiance phase boundary remains constant, independent of the coupling strengths.
+This suggests an unconventional SPT.
+In contrast, for other coupling phases, the phase boundary depends on both the cavity and atom-photon coupling strengths, indicative of a conventional SPT.
+We argue that the conventional SPT corresponds to a second-order quantum phase transition, whereas the unconventional SPT exhibits characteristics of a first-order quantum phase transition.
+Despite the model’s non-$\mathcal{PT}$-symmetry, we observe a fully real energy spectrum under certain conditions.
+Analysis of photon and atom populations, as well as fidelity measurements, indicates that the unconventional SPT lacks quantum criticality, characterized by the absence of asymptotic behavior and a sharp drop in fidelity at a single parameter value.
+In contrast, the conventional SPT exhibits quantum criticality, confirmed by critical exponents extracted from fidelity susceptibility scaling analyses, showing agreement with the single Rabi model’s universality class.
+Recent advancements in circuit-QED and ion trap experimental techniques offer substantial potential to implement and verify these identified SPTs in high-performance, state-of-the-art architectures.
+Moreover, these capabilities can be extended to investigate more complex coupled light-matter systems that currently surpass the capabilities of classical computational simulations.
 We acknowledge the support from NSFC under Grant No. 12174346.